### v1

# An Optomechanical Accelerometer Search for Ultralight Dark Matter

Abstract
[…]
Dark matter is one of the enduring scientific mysteries of our time. Astrophysical observations suggest that an unidentified particle or class of particles accounts for $\sim\!85\;\%$ of the Universe’s gravitating matter content; however, their specific mass and the nature of their coupling to Standard Model (SM) particles remains unknown, despite over 40 years of searching using a variety of direct and indirect techniques Bertone and Tait (2018).
Advances in cooling and probing of solid state mechanical oscillators using optical cavities Aspelmeyer et al. (2014) have spurred widespread interest in using such cavity optomechanical systems as tabletop dark matter (DM) detectors Carney et al. (2021a); Antypas et al. (2022), leveraging compatibility with cryogenics and rapidly maturing techniques that enable operation at fundamental quantum noise limits. Particle- and wave-like DM detectors have been proposed, based on impulsive Carney et al. (2020) and continuous Carney et al. (2021b) force measurement protocols, respectively. Massive arrays of optomechanical sensors might serve as DM track detectors Carney et al. (2020); Blanco et al. (2022) or phased antennae arrays Derevianko (2018), and have inspired investigation as an application of entanglement-enhanced distributed force sensing Brady et al. (2022, 2023).
In this Letter, we describe a search for wavelike, ultralight dark matter (UDM) using a cavity optomechanical system operating as an accelerometer. The concept of accelerometer-based UDM detection was introduced by Graham et al. (2016) and later reframed as an optomechanical force sensing problem by Carney et al. (2021b); Manley et al. (2021). The basic premise is that DM particles with mass $m_{\mathrm{DM}}\lesssim 10$ eV/c2 Jackson Kimball and Van Bibber (2023) (determined by the local dark matter energy density $\rho_{\mathrm{DM}}$) coherently combine to create an oscillating field, and that this field may act on SM atoms similar to the electromagnetic force—but with electric charge replaced with a generalized charge such as baryon minus lepton (B-L) number. As such, two free-falling SM objects exposed to UDM would experience a differential acceleration in proportion to $\sqrt{\rho_{\mathrm{DM}}}$ and their fractional charge difference $\Delta_{12}$ Manley et al. (2021); Carney et al. (2021b):
$a_{\mathrm{DM}}(t)=g\Delta_{\mathrm{12}}a_{0}\cos[\omega_{\mathrm{DM}}t+\phi_{\mathrm{DM}}(t)],$ (1)
where $g$ is an unknown coupling constant, $\omega_{\mathrm{DM}}$ is the UDM Compton frequency, $a_{0}=2.1\times 10^{11}\;\mathrm{m}/\mathrm{s}^{2}$ SI is a constant proportional to $\sqrt{\rho_{\mathrm{DM}}}$, and $\phi_{\mathrm{DM}}$ is a random phase accounting for spectral diffusion of DM particles due gravitational interaction with bodies in our Solar neighborhood.
Figure 1: Dual-membrane optomechanical dark matter detector. (a) Si chip with suspended Si3N4 membranes fixed to a Cu plate. Inset: Microscope image of trampoline (foreground) and square membrane (background) serving as acceleration test and reference mass, respectively. Concept: (b) UDM differentially accelerates the Cu plate, Si3N4 membranes, and Si chip. (c) In the reference frame of the base, the membranes experience a base excitation. Their different stiffnesses $k_{1,2}$ result in a relative displacement $x$.
Figure 2: Cryogenic operation enabled by vibration isolation. (a) Photos of Bluefors LD-4K cryostat (left) and custom vibration isolation system (VIS) based on thin Cu strips (right) suspended from the mixing stage. (b) Cavity alignment module (photo, bottom) and readout: probe and feedback laser beams are fiber-coupled and aligned to the dual-membrane cavity; the assembly is mounted on the VIS-platform (top, cartoon). A photodetector placed outside the cryostat records the cavity’s transmission. (c) Finite-element simulations of the fundamental modes of the VIS along $x$ (left) and $y$ (right). (d) Transmission fringes obtained from a laser detuning sweep before and after cooldown, showing cavity alignment remains stable. The green circle highlights the detuning used for side-of-fringe readout. (e) Energy ringdown of the trampoline’s fundamental mode at 4 K reveals $Q_{\mathrm{m}}$ of 60 million. (f) Broadband spectrum showing a reduction of vibration background above $f\sim 1\;\mathrm{Hz}$.
The design of our optomechanical accelerometer was proposed in Manley et al. (2021) and experimentally realized in Chowdhury et al. (2023), and consists of a pair of Si3N4 membranes with different stiffnesses—a trampoline and a square membrane—vertically integrated on a Si chip, forming a Fabry-Perot cavity. As illustrated in Fig. 1, fixing the cavity to a Cu plate (possessing a larger B-L [neutron] density than Si3N4 and Si) translates the UDM signal into an effective chip acceleration $a_{\mathrm{DM}}$. If coincident with the resonance frequency of the trampoline $\omega_{\mathrm{m}}$, the acceleration gets amplified by the mechanical quality factor into a displacement $x_{\mathrm{DM}}=Q_{\mathrm{m}}a_{\mathrm{DM}}/\omega_{\mathrm{m}}^{2}$ xDM , yielding sensitivity to UDM with a thermal-noise-equivalent coupling strength of Manley et al. (2021)
$g^{\mathrm{(th)}}\approx\sqrt{\frac{2k_{B}T}{mQ_{\mathrm{m}}Q_{\mathrm{DM}}}}\frac{\omega_{\mathrm{DM}}}{\Delta_{12}a_{0}}\times\left(\frac{2Q_{\mathrm{DM}}}{\omega_{\mathrm{DM}}\tau}\right)^{1/4}$ (2)
where $T$ is the bath temperature, $m$ is the effective mass of the trampoline (accounting for its modeshape and the finite stiffness of the square membrane Manley et al. (2021)), $Q_{\mathrm{DM}}\sim 10^{6}$ til is the effective $Q$ factor (number of coherent cycles) of the UDM field, and $\tau$ is the measurement time, assuming $\omega_{\mathrm{DM}}\tau>Q_{\mathrm{DM}}$.
Equation 2 illustrates the desirability of a high $m\times Q_{\mathrm{m}}$ mechanical oscillator operated at cryogenic temperatures for sustained intervals. To realize a novel sensitivity, Manley et al. (2021) proposed an $\omega_{\mathrm{m}}\sim 2\pi\times 1$ kHz, $Q_{\mathrm{m}}\sim 10^{9}$, cm-scale Si3N4 membrane ($m\sim 1\;\mathrm{mg}$) operating in a closed-cycle dilution refrigerator ($T\sim 10$ mK) for a year ($\tau\sim 10^{7}$ s), enabling $g_{\mathrm{DM}}^{\mathrm{(th)}}\sim 10^{-25}$—two orders of magnitude below current constraints set by precision torsion balance tests of the equivalence principle Wagner et al. (2012).
Figure 3: Displacement and acceleration measurements. (a) Calibrated membrane-trampoline displacement PSD near trampoline’s fundamental mode frequency $f_{\mathrm{m}}$, for a weak probe with transmitted power $P_{\mathrm{out}}\approx 1\;\upmu\mathrm{W}$. The red (gray) trace is obtained with (without) optical damping. (b) Corresponding closed-loop acceleration PSD after inverting the mechanical susceptibility. (c) Plots of acceleration noise versus frequency detuning from resonance at $P_{\mathrm{out}}=1\;\upmu\mathrm{W}$ (red) and $100\;\upmu\mathrm{W}$ (blue), illustrating the trade-off between bandwidth and sensitivity due to photothermal heating. The gray-shaded region corresponds to the UDM-signal linewidth $\gamma_{\mathrm{DM}}\approx 2\pi\times 0.08$ Hz. (d) Photothermal frequency tuning of $f_{\mathrm{m}}$. Increasing the probe power decreases $f_{\mathrm{m}}$ at the expense of increased thermal noise $S_{a}^{\mathrm{th}}$.
Here we report a more modest, first generation search for UDM employing a $\omega_{\mathrm{m}}\sim 2\pi\times 10$ kHz, $Q_{\mathrm{m}}\sim 10^{8}$, millimeter-scale Si3N4 trampoline ($m\sim 10\;\mathrm{ng}$) operating in a closed-cycle 4 K cryostat for $\tau\sim 10^{3}$ s, targeting $g_{\mathrm{DM}}^{\mathrm{(th)}}\sim 10^{-16}$. Our aim is to demonstrate the working principle behind a cavity optomechanical vector UDM detector, highlighting the challenges of vibration isolation, how to improve signal averaging using radiation pressure feedback cooling, strategies to extend bandwidth exploiting quantum-noise-limited readout and photothermal frequency tuning, and routines for data analysis by template matching to a model for the UDM signal.
An overview of our cryogenic apparatus and device characterization is shown in Fig. 2. In conceiving such a system, we first emphasize the challenging acceleration sensitivity requirements implied by current UDM constraints in the “optomechanics” band, $f_{\mathrm{DM}}=\omega_{\mathrm{DM}}/2\pi\sim 1\;\mathrm{Hz}$ to $1\;\mathrm{MHz}$. For B-L UDM, for example, $g\lesssim 10^{-22}$ (see Fig. 4) corresponds to an acceleration sensitivity of $\sqrt{S_{a}}\sim ga_{0}\Delta_{12}\sqrt{\omega_{\mathrm{DM}}/Q_{\mathrm{DM}}}\lesssim\Delta_{12}\times 10^{-11}g_{0}/\sqrt{\mathrm{Hz}}$, where $g_{0}=9.8\;\mathrm{m}/\mathrm{s}^{2}$ is Earth’s standard gravity. This sensitivity has been achieved by the Laser Interferomeric Gravitational Wave Observatory (LIGO) at frequencies below 1 kHz—indeed, B-L constraints from $100\;\mathrm{Hz}$ to $2\;\mathrm{kHz}$ currently belong to LIGO—however, at higher frequencies, we know of no such reported acceleration sensitivity, including with mechanical oscillators. Challenges include unfavorable thermal noise scaling $S_{a}^{\mathrm{th}}=4k_{B}T\omega_{\mathrm{m}}/(mQ_{\mathrm{m}})$, ambient vibrations produced by cryostats, and difficulty thermalizing micromechanical resonators to sufficiently low $T$, while at the same time being able to resolve their thermal motion.
We adopt a conservative approach and house our detector in a field-upgradable Bluefors LD-4K cryostat (LD with the dilution unit removed) NIS , using the extra sample volume for a two-foot-long (1 foot = 1 ft. = 30.5 cm) pendulum vibration isolation system (VIS) Planz et al. (2023) (Fig. 2a). The frequency of the pendulum, $f_{\mathrm{p}}=0.5$ Hz, is five orders of magnitude below the trampoline’s $f_{\mathrm{m}}=39$ kHz, implying access to an isolation factor of $(f_{\mathrm{m}}/f_{\mathrm{p}})^{2}\sim 10^{10}$. To mitigate pendulum swing, the accelerometer and fiber-optic delivery system are assembled and pre-aligned on a small optical bench mounted on the sample stage (Fig. 2b). The cavity transmission is monitored through a viewport by a low-noise Si photodetector. As shown in Fig. 2e, the cavity alignment survived cooldown with minimal loss of visibility. A mechanical ringdown revealed $Q=6\times 10^{7}$ and a set of thermal noise measurements (see below and SI ) confirmed that the device thermalized to the cryostat base temperature of 3.5 K, corresponding to $\sqrt{S_{a}^{\mathrm{th}}}=3\times 10^{-8}\;g_{0}/\sqrt{\mathrm{Hz}}$, using $m=12\;\mathrm{ng}$ Chowdhury et al. (2023).
To search for UDM, we analyze acceleration spectra for the presence of a narrow spectral feature with width $\gamma_{\mathrm{DM}}=\omega_{\mathrm{DM}}/Q_{\mathrm{DM}}$. Toward this end, Fig. 3 shows near-resonance estimates of the trampoline-membrane displacement power spectral density (PSD) $S_{x}[\omega]$ and, by inference, chip acceleration PSD $S_{a}[\omega]$, using a standard feedback-assisted force sensing protocol Chowdhury et al. (2023); Gavartin et al. (2012); Pontin et al. (2014); Gisler et al. (2024). For these measurements, the cavity was probed with a low noise wavelength-tunable (840-852 nm) diode laser, for displacement readout, and an auxiliary fixed wavelength (658 nm) diode laser, for radiation pressure feedback. The readout laser was tuned to the side of the fringe (Fig. 2e) to maximize sensitivity. The feedback laser was intensity modulated with a delayed copy of the readout photocurrent to damp the resonance by $\gamma>\gamma_{\mathrm{DM}}$ Col (Fig. 3a). Periodograms were then averaged for $\tau=800\;\mathrm{s}$ in $N=40$ intervals of $\tau_{\mathrm{p}}=20\;\mathrm{s}\approx 10\gamma_{\mathrm{DM}}^{-1}$, calibrated in displacement units using the fringe slope Chowdhury et al. (2023), and divided by the closed-loop acceleration susceptibility $\chi[\omega]=(\omega^{2}-\omega_{\mathrm{m}}^{2}+i\gamma\omega_{\mathrm{m}})^{-1}$, yielding an acceleration PSD estimate $\hat{S}_{a}[\omega]$ with an ideal (Gaussian-distributed) mean and standard deviation of
$S_{a}[\omega]$ $=S_{a}^{\mathrm{DM}}[\omega]+S_{a}^{\mathrm{th}}+|\chi[\omega]|^{-2}S_{x}^{\mathrm{imp}}\;\mathrm{and}$ (3a)
$\sigma_{\hat{S}_{a}}[\omega]$ $=S_{a}[\omega]/\sqrt{N}$ (3b)
respectively, where $S_{x}^{\mathrm{imp}}$ is the displacement readout noise (imprecision), and $S_{a}^{\mathrm{DM}}[\omega]$ is the hypothetical UDM signal.
As shown in Fig. 3, for readout powers from $P_{\mathrm{out}}\approx 1\;\upmu\mathrm{W}$ to $100\;\upmu\mathrm{W}$, we realize a shot-noise-limited displacement imprecision of $\sqrt{S_{x}^{\text{im\smash{p}}}}\approx(10^{-13}\text{--}10^{-14})\;\mathrm{m}/\sqrt{\mathrm{Hz}}$, corresponding to a thermal-noise-limited bandwidth of $\Delta\omega_{\mathrm{th}}=\omega_{\mathrm{m}}\sqrt{S_{x}^{\mathrm{th}}/S_{x}^{\text{im\smash{p}}}}\approx 2\pi\times 2\;\mathrm{Hz}$, or $\Delta\omega_{\mathrm{th}}/\gamma_{\mathrm{DM}}\approx 26$ independent DM bins. We also observe photothermal heating at the level of $dT/dP_{\mathrm{out}}\approx 0.4\;\mathrm{K}/\upmu\mathrm{W}$. While anomalously large, this heating enables photothermal tuning of the resonance frequency over 25 Hz (330 DM bins), at the expense of a 3-fold increase in $\sqrt{S_{a}^{\mathrm{th}}}$.
Our search algorithm SI involves matched-filtering $\hat{S}_{a}[\omega]$ to a model $S_{a}^{\mathrm{DM}}[\omega]=\langle a_{\mathrm{DM}}^{2}\rangle G_{\mathrm{DM}}[\omega]$, where $\langle a_{\mathrm{DM}}^{2}\rangle$ is the unknown UDM acceleration power and $G_{\mathrm{DM}}$ is the normalized lineshape ($\int G_{\mathrm{DM}}[\omega]d\omega/(2\pi)=1$). Towards this end, we consider an astrophysically motivated lineshape Derevianko (2018); Antypas et al. (2022)
$G_{\mathrm{DM}}[\omega]\approx\frac{\sqrt{8\pi}}{e\gamma_{\mathrm{DM}}}e^{-\tfrac{2(\omega-\omega_{\mathrm{DM}})}{\gamma_{\mathrm{DM}}}}\sinh\sqrt{1+\tfrac{4(\omega-\omega_{\mathrm{DM}})}{\gamma_{\mathrm{DM}}}}$ (4)
[…]
$\hat{D}[\omega_{\mathrm{DM}}]=\sigma^{2}_{\hat{D}}[\omega_{\mathrm{DM}}]\sum_{ik}\frac{G_{\mathrm{DM}}^{\tau_{\mathrm{p}}}[\omega_{k},\omega_{\mathrm{DM}}]}{\sigma^{2}_{\hat{S}^{(i)}_{a}}[\omega_{k}]}\left(\hat{S}^{\smash{(i)}}_{a}[\omega_{k}]-S_{a}^{\mathrm{n},(i)}[\omega_{k}]\right)$ (5)
where $G_{\mathrm{DM}}^{\tau_{\mathrm{p}}}$ is UDM lineshape binned over $\tau_{\mathrm{p}}^{-1}$ and $S_{a}^{\mathrm{n}}=S_{a}-S_{a}^{\mathrm{DM}}$ is the total measurement noise. Applying frequentist and Bayesian frameworks Fre , respectively, we implicitly define a detection threshold $D_{\mathrm{DT}}$ and an upper bound for the DM signal power $D_{\mathrm{UL}}$, based on a confidence level
$\mathrm{CL}=\left(\int_{-\infty}^{D_{\mathrm{DT}}}\rho(\hat{D}|D=0)d\hat{D}\right)^{N_{\mathrm{b}}}=\int_{-\infty}^{D_{\mathrm{UL}}}\rho(D|\hat{D})dD,$ (6)
[…]
Figure 4: Search for B-L UDM. (a) Estimates of UDM acceleration power $\langle a_{\mathrm{DM}}^{2}\rangle\equiv D$ obtained from the data in Fig. 3d using a matched-filter and weighted averaging, Eq. 5. Dashed lines are thresholds for confidence level $\mathrm{CL}=68\;\%$ ($1\sigma$) and $95\;\%$ ($2\sigma$). The solid red line is our threshold, which includes the look-elsewhere effect (Eq. 6). (b) Histogram of normalized power estimates $\hat{D}/\sigma_{D}$. Solid and dotted curves show measured and ideal distributions expected from periodogram averaging (Eq. 3b). (c) Constraints on $g_{\mathrm{B-L}}$ from data in (a). Light and bold traces are statistical and analytic estimates, respectively. Gray traces assume a rigid chip; red traces include the estimated chip transfer function SI . (d) Comparison to current constraints from LIGO Guo et al. (2019); Abbott et al. (2022), the Eöt-Wash experiment Wagner et al. (2012), MICROSCOPE Bergé et al. (2018), and POLONAISE Amaral et al. (2025). Gray dashed line is an extension of the analytical model in (c) assuming a displacement imprecision of $S_{x}^{\mathrm{imp}}=(5\times 10^{-14}\;\mathrm{m}/\sqrt{\mathrm{Hz}})^{2}$ SI .
Results of the UDM search are summarized in Fig. 5. As shown in Fig. 5(a,b), we observe a near-Gaussian distribution of weighted estimators $\hat{D}[\omega_{k}]/\sigma_{\hat{D}}[\omega_{k}]$, with variance $\sigma^{2}_{\hat{D}}[\omega_{k}]$ SI minimized over the $25$ Hz resonant scan window in Fig. 3d. This motivates a likelihood function $\rho(\hat{D}|D)\propto\mathrm{exp}[(D-\hat{D})^{2}/2\sigma^{2}_{\hat{D}}]$, from which we determine the $\mathrm{CL}=95\;\%$ confidence threshold shown in Fig. 5a. Accounting for the look-elsewhere effect over a search span of $\Delta f=1\;\mathrm{kHz}$ ($N_{\mathrm{b}}=\tau_{\mathrm{p}}\Delta f=10^{4}$), we observe no detection events.
To set a limit on the DM coupling strength, we combine the experimentally determined likelihood function with Bayes theorem and a prior $\rho(D)$ that is uniform for $D>0$, yielding a median limit $D_{\mathrm{UL}}\approx\sqrt{2}\sigma_{\hat{D}}\mathrm{Erf}^{-1}[\mathrm{CL}]$ SI and a $2\sigma$ bound
$g\leq\frac{\sqrt{2D_{\mathrm{CL}}}}{a_{0}\Delta_{12}}\approx\frac{2\sqrt{\sigma_{\hat{D}}}}{a_{0}\Delta_{12}}$ (7)
for $\mathrm{CL}=95\;\%$. We consider B-L coupling ($g_{\mathrm{B-L}}$) as a canonical example (reflecting its origin as a minimal, anomaly-free gauge extension of the SM Fileviez Pérez and Wise (2010); Bauer et al. (2018)), in which case $\Delta_{12}=Z_{1}/A_{1}-Z_{2}/A_{2}$, where $Z_{i}$ and $A_{i}$ are the atomic and mass number of body $i$, respectively. For Si3N4 and Si, $Z/A\approx 0.50$ and for Cu, $Z/A\approx 0.46$, yielding $\Delta_{12}\approx 0.04$. Using this value presumes that the accelerometer behaves like a heterogeneous mechanical dimer (two bodies made of different material, attached by a simple spring) Manley et al. (2021), as substantiated in SI .
Inferred $g_{\mathrm{B-L}}$ constraints are shown in Fig. 5c, assuming a perfectly rigid chip (gray) and correcting for the estimated chip response function (red) SI . The lower bound $g_{\mathrm{B-L}}\approx 1\times 10^{-17}$ corresponds to an acceleration amplitude resolution of $\sqrt{\sigma_{\hat{D}}}\approx 4\times 10^{-9}\;g_{0}$ and PSD resolution of $\sigma_{\hat{S}_{a}}(\omega_{\mathrm{DM}})\approx(2\times 10^{-8}g_{0}/\sqrt{\mathrm{Hz}})^{2}$, consistent with the variance of the lowest PSD estimate in Fig. 3 SI (at $T\approx$ 8 K due to photothermal heating). Shown in Fig. 5d is our bound combined with the best current constraints due the Eöt-Wash experiment Wagner et al. (2012) and various others at lower frequency.
Looking forward, realizing the originally proposed metrics Manley et al. (2021) requires operating at 10 mK, increasing the cavity finesse to $\mathcal{F}=100$, and increasing the size and $Q$ of the membrane to approximately $10$ cm and $10^{9}$, respectively. In principle this is possible: Sub-100-mK, $\mathcal{F}\gtrsim 10^{4}$ cavity optomechanics experiments with Si3N4 membranes have been demonstrated by multiple groups, including a recent study Planz et al. (2023) employing a pendulum VIS on which ours is based. High reflectivity photonic crystal (PtC) membranes have also been demonstrated, including stable $\mathcal{F}\approx 10^{3}$ cavities employing a gradient-pitch PtC Agrawal et al. (2024). In terms of size, decimeter-scale membranes are now commercially available Nor , and centimeter-scale PtC membranes Moura et al. (2018); Norder et al. (2025) have been realized, driven by the Starshot project. Finally, the expected mechanical $Q$ versus size scaling (due to dissipation dilution) from $10^{8}$ to $10^{10}$ has been verified with a centimeter-scale Si3N4 beam Cupertino et al. (2024).
An alternative approach to larger membranes is 3D-mass-loading Pratt et al. (2023); Zhou et al. (2021); Bawden et al. (2025). Towards this end, we highlight a recent demonstration of $m\sim 0.1\;\mathrm{mg}$, $f_{\mathrm{m}}\sim 100$ Hz, $Q_{\mathrm{m}}\sim 10^{7}$ torsion micropendula based on mass-loaded Si3N4 nanoribbons, which achieve $S_{a}^{\mathrm{th}}\sim 10^{-9}\;g_{0}/\sqrt{\mathrm{Hz}}$ at room temperature Condos et al. (2025). Cryogenically cooled, these devices could reach the desired sensitivity for novel UDM searches in the 10 Hz to 1 kHz band, in conjunction with quantum-limited optical lever measurements Pluchar et al. (2025); Shin et al. (2025); they also provide a route to vibration isolation through heterogeneous mass-loading Sun et al. (2025).
Finally, we highlight a recent proposal to search for UDM with an array of $N=10$ cryogenic membrane accelerometers probed by a distributed-squeezed light source Brady et al. (2023), building on an $N=2$ demonstration with two $0.1$-$m$, $f_{\mathrm{m}}=6\;\mathrm{MHz}$ membranes Xia et al. (2023). This approach enables enhanced sensitivity (as much as $N$-fold Brady et al. (2023)) to a distributed force if the power per sensor is constrained and, therefore, a route towards improved UDM searches with a cryogenic detector limited by photothermal heating—in addition to common mode noise rejection and bandwidth enhancement. Loss and scalability are key challenges. Towards this end, an emerging class of photonic-integrated optomechanical accelerometers Ge et al. (2025) and squeezed light sources Clark et al. (2025) offer a promising route.
[…]
I Acknowledgments
The authors thank Swati Singh and Daniel Grin for laying the theoretical foundations of this work, Atkin Hyatt for designing and fabricating next generation test masses presented in the Supporting Materials, Christian Pluchar for guidance with laser intensity noise reduction, Andrew Land for help in machining and assembling the vibration isolation system, and Utkal Pandurangi for helpful discussions. This work was supported by NSF Grants 2209473 and 1945832.