### v2

# Optomechanical Accelerometer Search for Ultralight Dark Matter

Abstract
Cavity optomechanical systems have recently been proposed as detectors for ultralight dark matter, leveraging their ability to cool and probe mechanical oscillators at the quantum limit. Here we present a resonant search for ultralight dark matter using a cavity optomechanical accelerometer. The detector consists of a cryogenic Si3N4-membrane cavity mounted to a 4 K copper plate, with photothermal tuning used to scan its 39 kHz mechanical resonance. Shot-noise-limited displacement readout and radiation-pressure feedback cooling yield an acceleration sensitivity of 10 $\mathrm{n}g_{0}/\sqrt{\mathrm{Hz}}$ over 30 Hz near resonance. The detector’s material inhomogeneity gives access to direct vector coupling to the dark-matter field. We conduct a Bayesian search based on matched-filter statistics, yielding upper bounds consistent with thermal noise and above those set by equivalence principle tests. No signal is observed, but the experiment demonstrates stable, quantum-limited operation and validates a scalable approach to resonant detection. With optimized test masses, lower temperature, and multiplexed arrays, the platform offers a path toward competitive constraints on vector-mediated dark-matter interactions.
Dark matter is one of the enduring scientific mysteries of our time. Astrophysical observations suggest that an unidentified particle or class of particles accounts for $\sim\!85\;\%$ of the Universe’s gravitating matter content; however, their specific mass and the nature of their coupling to Standard Model (SM) particles remains unknown, despite over 40 years of searching using a variety of direct and indirect techniques [1].
Advances in cooling and probing of solid state mechanical oscillators using optical cavities [2] have spurred widespread interest in using such cavity optomechanical systems as tabletop dark matter (DM) detectors [3, 4, 5, 6, 7, 8, 9, 10], leveraging compatibility with cryogenics and rapidly maturing techniques that enable operation at fundamental quantum noise limits. Particle- and wave-like DM detectors have been proposed, based on impulsive [11] and continuous [12] force measurement protocols, respectively. Massive arrays of optomechanical sensors might serve as DM track detectors [11, 13] or phased antennae arrays [14], and have inspired investigation as an application of entanglement-enhanced distributed force sensing [15, 16].
In this Letter, we describe a search for wavelike, ultralight dark matter (UDM) using a cavity optomechanical system operating as an accelerometer. The concept of accelerometer-based UDM detection was introduced by [17] and later reframed as an optomechanical force sensing problem by [12, 7]. The basic premise is that DM particles with mass $m_{\mathrm{DM}}\lesssim 10$ eV/c2 [18] (determined by the local dark matter energy density $\rho_{\mathrm{DM}}$) coherently combine to create an oscillating field, and that this field may act on SM atoms similar to the electromagnetic force—but with electric charge replaced with a generalized charge such as baryon minus lepton (B-L) number. As such, two free-falling SM objects exposed to UDM would experience a differential acceleration in proportion to $\sqrt{\rho_{\mathrm{DM}}}$ and their fractional charge difference $\Delta_{12}$ [7, 12]:
$a_{\mathrm{DM}}(t)=g\Delta_{\mathrm{12}}a_{0}\cos[\omega_{\mathrm{DM}}t+\phi_{\mathrm{DM}}(t)],$ (1)
where $g$ is an unknown coupling constant, $\omega_{\mathrm{DM}}$ is the UDM Compton frequency, $a_{0}=2.1\times 10^{11}\;\mathrm{m}/\mathrm{s}^{2}$ [19] is a constant proportional to $\sqrt{\rho_{\mathrm{DM}}}$, and $\phi_{\mathrm{DM}}$ is a random phase accounting for spectral diffusion of DM particles due gravitational interaction with bodies in our Solar neighborhood.
Figure 1: Dual-membrane optomechanical dark matter detector. (a) Si chip with suspended Si3N4 membranes fixed to a Cu plate. Inset: Microscope image of trampoline (foreground) and square membrane (background) serving as acceleration test and reference mass, respectively. Concept: (b) UDM differentially accelerates the Cu plate, Si3N4 membranes, and Si chip. (c) In the reference frame of the base, the membranes experience a base excitation. Their different stiffnesses $k_{1,2}$ result in a relative displacement $x$.
Figure 2: Cryogenic operation enabled by vibration isolation. (a) Photos of Bluefors LD-4K cryostat (left) and custom vibration isolation system (VIS) based on thin Cu strips (right) suspended from the mixing stage. (b) Cavity alignment module (photo, bottom) and readout: probe and feedback laser beams are fiber-coupled and aligned to the dual-membrane cavity; the assembly is mounted on the VIS-platform (top, cartoon). A photodetector placed outside the cryostat records the cavity’s transmission. (c) Finite-element simulations of the fundamental modes of the VIS along $x$ (left) and $y$ (right). (d) Transmission fringes obtained from a laser detuning sweep before and after cooldown, showing cavity alignment remains stable. The green circle highlights the detuning used for side-of-fringe readout. (e) Energy ringdown of the trampoline’s fundamental mode at 4 K reveals $Q_{\mathrm{m}}$ of 60 million. (f) Broadband spectrum showing a reduction of vibration background above $f\sim 1\;\mathrm{Hz}$.
The design of our optomechanical accelerometer was proposed in [7] and experimentally realized in [20], and consists of a pair of Si3N4 membranes with different stiffnesses—a trampoline and a square membrane—vertically integrated on a Si chip, forming a Fabry-Perot cavity. As illustrated in Fig. 1, fixing the cavity to a Cu plate (possessing a larger B-L [neutron] density than Si3N4 and Si) translates the UDM signal into an effective chip acceleration $a_{\mathrm{DM}}$. If coincident with the resonance frequency of the trampoline $\omega_{\mathrm{m}}$, the acceleration gets amplified by the mechanical quality factor into a displacement $x_{\mathrm{DM}}=Q_{\mathrm{m}}a_{\mathrm{DM}}/\omega_{\mathrm{m}}^{2}$ [21], yielding sensitivity to UDM with a thermal-noise-equivalent coupling strength of [7]
$g^{\mathrm{(th)}}\approx\sqrt{\frac{2k_{B}T}{mQ_{\mathrm{m}}Q_{\mathrm{DM}}}}\frac{\omega_{\mathrm{DM}}}{\Delta_{12}a_{0}}\times\left(\frac{2Q_{\mathrm{DM}}}{\omega_{\mathrm{DM}}\tau}\right)^{1/4}$ (2)
where $T$ is the bath temperature, $m$ is the effective mass of the trampoline (accounting for its modeshape and the finite stiffness of the square membrane [7]), $Q_{\mathrm{DM}}\sim 10^{6}$ [22] is the effective $Q$ factor (number of coherent cycles) of the UDM field, and $\tau$ is the measurement time, assuming $\omega_{\mathrm{DM}}\tau>Q_{\mathrm{DM}}$.
Equation 2 illustrates the desirability of a high $m\times Q_{\mathrm{m}}$ mechanical oscillator operated at cryogenic temperatures for sustained intervals. To realize a novel sensitivity, [7] proposed an $\omega_{\mathrm{m}}\sim 2\pi\times 1$ kHz, $Q_{\mathrm{m}}\sim 10^{9}$, cm-scale Si3N4 membrane ($m\sim 1\;\mathrm{mg}$) operating in a closed-cycle dilution refrigerator ($T\sim 10$ mK) for a year ($\tau\sim 10^{7}$ s), enabling $g_{\mathrm{DM}}^{\mathrm{(th)}}\sim 10^{-25}$—two orders of magnitude below current constraints set by precision torsion balance tests of the equivalence principle [23, 24].
Figure 3: Displacement and acceleration measurements. (a) Calibrated membrane-trampoline displacement PSD near trampoline’s fundamental mode frequency $f_{\mathrm{m}}$, for a weak probe with transmitted power $P_{\mathrm{out}}\approx 1\;\upmu\mathrm{W}$. The red (gray) trace is obtained with (without) optical damping. (b) Corresponding closed-loop acceleration PSD after inverting the mechanical susceptibility. (c) Plots of acceleration noise versus frequency detuning from resonance at $P_{\mathrm{out}}=1\;\upmu\mathrm{W}$ (red) and $100\;\upmu\mathrm{W}$ (blue), illustrating the trade-off between bandwidth and sensitivity due to photothermal heating. The gray-shaded region corresponds to the UDM-signal linewidth $\gamma_{\mathrm{DM}}\approx 2\pi\times 0.08$ Hz. (d) Photothermal frequency tuning of $f_{\mathrm{m}}$. Increasing the probe power decreases $f_{\mathrm{m}}$ at the expense of increased thermal noise $S_{a}^{\mathrm{th}}$.
Here we report a more modest, first generation search for UDM employing a $\omega_{\mathrm{m}}\sim 2\pi\times 10$ kHz, $Q_{\mathrm{m}}\sim 10^{8}$, millimeter-scale Si3N4 trampoline ($m\sim 10\;\mathrm{ng}$) operating in a closed-cycle 4 K cryostat for $\tau\sim 10^{3}$ s, targeting $g_{\mathrm{DM}}^{\mathrm{(th)}}\sim 10^{-16}$. Our aim is to demonstrate the working principle behind a cavity optomechanical vector UDM detector, highlighting the challenges of vibration isolation, how to improve signal averaging using radiation pressure feedback cooling, strategies to extend bandwidth exploiting quantum-noise-limited readout and photothermal frequency tuning, and routines for data analysis by template matching to a model for the UDM signal.
An overview of our cryogenic apparatus and device characterization is shown in Fig. 2. In conceiving such a system, we first emphasize the challenging acceleration sensitivity requirements implied by current UDM constraints in the “optomechanics” band, $f_{\mathrm{DM}}=\omega_{\mathrm{DM}}/2\pi\sim 1\;\mathrm{Hz}$ to $1\;\mathrm{MHz}$. For B-L UDM, for example, $g\lesssim 10^{-22}$ (see Fig. 4) corresponds to an acceleration sensitivity of $\sqrt{S_{a}}\sim ga_{0}\Delta_{12}\sqrt{\omega_{\mathrm{DM}}/Q_{\mathrm{DM}}}\lesssim\Delta_{12}\times 10^{-11}g_{0}/\sqrt{\mathrm{Hz}}$, where $g_{0}=9.8\;\mathrm{m}/\mathrm{s}^{2}$ is Earth’s standard gravity. This sensitivity has been achieved by the Laser Interferomeric Gravitational Wave Observatory (LIGO) at frequencies below 1 kHz—indeed, B-L constraints from $100\;\mathrm{Hz}$ to $2\;\mathrm{kHz}$ currently belong to LIGO—however, at higher frequencies, we know of no such reported acceleration sensitivity, including with mechanical oscillators. Challenges include unfavorable thermal noise scaling $S_{a}^{\mathrm{th}}=4k_{B}T\omega_{\mathrm{m}}/(mQ_{\mathrm{m}})$, ambient vibrations produced by cryostats, and difficulty thermalizing micromechanical resonators to sufficiently low $T$, while at the same time being able to resolve their thermal motion.
We adopt a conservative approach and house our detector in a field-upgradable Bluefors LD-4K cryostat (LD with the dilution unit removed) [25], using the extra sample volume for a two-foot-long (1 foot = 1 ft. = 30.5 cm) pendulum vibration isolation system (VIS) [26] (Fig. 2a). The frequency of the pendulum, $f_{\mathrm{p}}=0.5$ Hz, is five orders of magnitude below the trampoline’s $f_{\mathrm{m}}=39$ kHz, implying access to an isolation factor of $(f_{\mathrm{m}}/f_{\mathrm{p}})^{2}\sim 10^{10}$. To mitigate pendulum swing, the accelerometer and fiber-optic delivery system are assembled and pre-aligned on a small optical bench mounted on the sample stage (Fig. 2b). The cavity transmission is monitored through a viewport by a low-noise Si photodetector. As shown in Fig. 2e, the cavity alignment survived cooldown with minimal loss of visibility. A mechanical ringdown revealed $Q=6\times 10^{7}$ and a set of thermal noise measurements (see below and [19]) confirmed that the device thermalized to the cryostat base temperature of 3.5 K, corresponding to $\sqrt{S_{a}^{\mathrm{th}}}=3\times 10^{-8}\;g_{0}/\sqrt{\mathrm{Hz}}$, using $m=12\;\mathrm{ng}$ [20].
To search for UDM, we analyze acceleration spectra for the presence of a narrow spectral feature with width $\gamma_{\mathrm{DM}}=\omega_{\mathrm{DM}}/Q_{\mathrm{DM}}$. Toward this end, Fig. 3 shows near-resonance estimates of the trampoline-membrane displacement power spectral density (PSD) $S_{x}[\omega]$ and, by inference, chip acceleration PSD $S_{a}[\omega]$, using a standard feedback-assisted force sensing protocol [20, 27, 28, 29]. For these measurements, the cavity was probed with a low noise wavelength-tunable (840-852 nm) diode laser, for displacement readout, and an auxiliary fixed wavelength (658 nm) diode laser, for radiation pressure feedback. The readout laser was tuned to the side of the fringe (Fig. 2e) to maximize sensitivity. The feedback laser was intensity modulated with a delayed copy of the readout photocurrent to damp the resonance by $\gamma>\gamma_{\mathrm{DM}}$ [30] (Fig. 3a). Periodograms were then averaged for $\tau=800\;\mathrm{s}$ in $N=40$ intervals of $\tau_{\mathrm{p}}=20\;\mathrm{s}\approx 10\gamma_{\mathrm{DM}}^{-1}$, calibrated in displacement units using the fringe slope [20], and divided by the closed-loop acceleration susceptibility $\chi[\omega]=(\omega^{2}-\omega_{\mathrm{m}}^{2}+i\gamma\omega_{\mathrm{m}})^{-1}$, yielding an acceleration PSD estimate $\hat{S}_{a}[\omega]$ with an ideal (Gaussian-distributed) mean and standard deviation of
$S_{a}[\omega]$ $=S_{a}^{\mathrm{DM}}[\omega]+S_{a}^{\mathrm{th}}+|\chi[\omega]|^{-2}S_{x}^{\mathrm{imp}}\;\mathrm{and}$ (3a)
$\sigma_{\hat{S}_{a}}[\omega]$ $=S_{a}[\omega]/\sqrt{N}$ (3b)
respectively, where $S_{x}^{\mathrm{imp}}$ is the displacement readout noise (imprecision), and $S_{a}^{\mathrm{DM}}[\omega]$ is the hypothetical UDM signal.
As shown in Fig. 3, for readout powers from $P_{\mathrm{out}}\approx 1\;\upmu\mathrm{W}$ to $100\;\upmu\mathrm{W}$, we realize a shot-noise-limited displacement imprecision of $\sqrt{S_{x}^{\text{im\smash{p}}}}\approx(10^{-13}\text{--}10^{-14})\;\mathrm{m}/\sqrt{\mathrm{Hz}}$, corresponding to a thermal-noise-limited bandwidth of $\Delta\omega_{\mathrm{th}}=\omega_{\mathrm{m}}\sqrt{S_{x}^{\mathrm{th}}/S_{x}^{\text{im\smash{p}}}}\approx 2\pi\times 2\;\mathrm{Hz}$, or $\Delta\omega_{\mathrm{th}}/\gamma_{\mathrm{DM}}\approx 26$ independent DM bins. We also observe photothermal heating at the level of $dT/dP_{\mathrm{out}}\approx 0.4\;\mathrm{K}/\upmu\mathrm{W}$. While anomalously large, this heating enables photothermal tuning of the resonance frequency over 25 Hz (330 DM bins), at the expense of a 3-fold increase in $\sqrt{S_{a}^{\mathrm{th}}}$.
Our search algorithm [19] involves matched-filtering $\hat{S}_{a}[\omega]$ to a model $S_{a}^{\mathrm{DM}}[\omega]=\langle a_{\mathrm{DM}}^{2}\rangle G_{\mathrm{DM}}[\omega]$, where $\langle a_{\mathrm{DM}}^{2}\rangle$ is the unknown UDM acceleration power and $G_{\mathrm{DM}}$ is the normalized lineshape ($\int G_{\mathrm{DM}}[\omega]d\omega/(2\pi)=1$). Towards this end, we consider an astrophysically motivated lineshape [14, 4]
$G_{\mathrm{DM}}[\omega]\approx\frac{\sqrt{8\pi}}{e\gamma_{\mathrm{DM}}}e^{-\tfrac{2(\omega-\omega_{\mathrm{DM}})}{\gamma_{\mathrm{DM}}}}\sinh\sqrt{1+\tfrac{4(\omega-\omega_{\mathrm{DM}})}{\gamma_{\mathrm{DM}}}}$ (4)
[…]
$\hat{D}[\omega_{\mathrm{DM}}]=\sigma^{2}_{\hat{D}}[\omega_{\mathrm{DM}}]\sum_{ik}\frac{G_{\mathrm{DM}}^{\tau_{\mathrm{p}}}[\omega_{k},\omega_{\mathrm{DM}}]}{\sigma^{2}_{\hat{S}^{(i)}_{a}}[\omega_{k}]}\left(\hat{S}^{\smash{(i)}}_{a}[\omega_{k}]-S_{a}^{\mathrm{n},(i)}[\omega_{k}]\right)$ (5)
where $G_{\mathrm{DM}}^{\tau_{\mathrm{p}}}$ is UDM lineshape binned over $\tau_{\mathrm{p}}^{-1}$ and $S_{a}^{\mathrm{n}}=S_{a}-S_{a}^{\mathrm{DM}}$ is the total measurement noise. Applying frequentist and Bayesian frameworks [31], respectively, we implicitly define a detection threshold $D_{\mathrm{DT}}$ and an upper bound for the DM signal power $D_{\mathrm{UL}}$, based on a confidence level
$\mathrm{CL}=\left(\int_{-\infty}^{D_{\mathrm{DT}}}\rho(\hat{D}|D=0)d\hat{D}\right)^{N_{\mathrm{b}}}=\int_{-\infty}^{D_{\mathrm{UL}}}\rho(D|\hat{D})dD,$ (6)
[…]
Figure 4: Search for B-L UDM. (a) Estimates of UDM acceleration power $\langle a_{\mathrm{DM}}^{2}\rangle\equiv D$ obtained from the data in Fig. 3d using a matched-filter and weighted averaging, Eq. 5. Dashed lines are thresholds for confidence level $\mathrm{CL}=68\;\%$ ($1\sigma$) and $95\;\%$ ($2\sigma$). The solid red line is our threshold, which includes the look-elsewhere effect (Eq. 6). (b) Histogram of normalized power estimates $\hat{D}/\sigma_{D}$. Solid and dotted curves show measured and ideal distributions expected from periodogram averaging (Eq. 3b). (c) Constraints on $g_{\mathrm{B-L}}$ from data in (a). Light and bold traces are statistical and analytic estimates, respectively. Gray traces assume a rigid chip; red traces include the estimated chip transfer function [19]. (d) Comparison to current constraints from LIGO [32, 33], the Eöt-Wash experiment [23], MICROSCOPE [34], and POLONAISE [35]. Gray dashed line is an extension of the analytical model in (c) assuming a displacement imprecision of $S_{x}^{\mathrm{imp}}=(5\times 10^{-14}\;\mathrm{m}/\sqrt{\mathrm{Hz}})^{2}$ [19].
Results of the UDM search are summarized in Fig. 4. As shown in Fig. 4(a,b), we observe a near-Gaussian distribution of weighted estimators $\hat{D}[\omega_{k}]/\sigma_{\hat{D}}[\omega_{k}]$, with variance $\sigma^{2}_{\hat{D}}[\omega_{k}]$ [19] minimized over the $25$ Hz resonant scan window in Fig. 3d. This motivates a likelihood function $\rho(\hat{D}|D)\propto\mathrm{exp}[(D-\hat{D})^{2}/2\sigma^{2}_{\hat{D}}]$, from which we determine the $\mathrm{CL}=95\;\%$ confidence threshold shown in Fig. 4a. Accounting for the look-elsewhere effect over a search span of $\Delta f=1\;\mathrm{kHz}$ ($N_{\mathrm{b}}=\tau_{\mathrm{p}}\Delta f=10^{4}$), we observe no detection events.
To set a limit on the DM coupling strength, we combine the experimentally determined likelihood function with Bayes theorem and a prior $\rho(D)$ that is uniform for $D>0$, yielding a median limit $D_{\mathrm{UL}}\approx\sqrt{2}\sigma_{\hat{D}}\mathrm{Erf}^{-1}[\mathrm{CL}]$ [19] and a $2\sigma$ bound
$g\leq\frac{\sqrt{2D_{\mathrm{CL}}}}{a_{0}\Delta_{12}}\approx\frac{2\sqrt{\sigma_{\hat{D}}}}{a_{0}\Delta_{12}}$ (7)
for $\mathrm{CL}=95\;\%$. We consider B-L coupling ($g_{\mathrm{B-L}}$) as a canonical example (reflecting its origin as a minimal, anomaly-free gauge extension of the SM [36, 37]), in which case $\Delta_{12}=Z_{1}/A_{1}-Z_{2}/A_{2}$, where $Z_{i}$ and $A_{i}$ are the atomic and mass number of body $i$, respectively. For Si3N4 and Si, $Z/A\approx 0.50$ and for Cu, $Z/A\approx 0.46$, yielding $\Delta_{12}\approx 0.04$. Using this value presumes that the accelerometer behaves like a heterogeneous mechanical dimer (two bodies made of different material, attached by a simple spring) [7], as substantiated in [19].
Inferred $g_{\mathrm{B-L}}$ constraints are shown in Fig. 4c, assuming a perfectly rigid chip (gray) and correcting for the estimated chip response function (red) [19]. The lower bound $g_{\mathrm{B-L}}\approx 1\times 10^{-17}$ corresponds to an acceleration amplitude resolution of $\sqrt{\sigma_{\hat{D}}}\approx 4\times 10^{-9}\;g_{0}$ and PSD resolution of $\sigma_{\hat{S}_{a}}(\omega_{\mathrm{DM}})\approx(2\times 10^{-8}g_{0}/\sqrt{\mathrm{Hz}})^{2}$, consistent with the variance of the lowest PSD estimate in Fig. 3 [19] (at $T\approx$ 8 K due to photothermal heating). Shown in Fig. 4d is our bound combined with the best current constraints from the Eöt-Wash experiment [23, 24] and various others at lower frequency.
Looking forward, realizing the original design sensitivity $g_{\mathrm{B-L}}\sim 10^{-25}$ [7] requires operating at 10 mK, increasing the cavity finesse to $\mathcal{F}=100$, increasing the size and $Q$ of the membrane to $10$ cm and $10^{9}$, respectively, vibration isolation at the $10^{-12}g_{0}/\sqrt{\mathrm{Hz}}$ level, and integrating for a year. Recent advances augur well for these upgrades: Sub-100-mK, $\mathcal{F}\gtrsim 10^{4}$ cavity optomechanics experiments with Si3N4 membranes have been demonstrated by several groups [26]; as have high reflectivity photonic crystal (PtC) membranes, including stable $\mathcal{F}\approx 10^{3}$ cavities employing a gradient-pitch PtC [38]. Decimeter-scale membranes are now commercially available [39], and centimeter-scale PtC membranes [40, 41] have been realized, driven by the Starshot project. The expected mechanical $Q$ versus size scaling (due to dissipation dilution [42]) from $10^{8}$ to $10^{10}$ has also been confirmed with centimeter-scale Si3N4 beams [43]. Finally, low vibration, closed-cycle (continuous operation) dilution refrigerator technology continues to advance in response to the superconducting quantum computing industry, with custom $10^{-11}g_{0}/\sqrt{\mathrm{Hz}}$ VIS demonstrated for frequencies as low as $10$ Hz [35].
A key component to realizing novel contraints will be optimizing the membrane geometry. In-plane patterning can enable ultrahigh-$Q$ [42], however, a promising alternative is 3D-mass-loading [44, 45, 46], which can enhance the more relevant $Q$-$m$ product (Eq. 2). Towards this end, we highlight a recent demonstration of $m\sim 0.1\;\mathrm{mg}$, $f_{\mathrm{m}}\sim 100$ Hz, $Q_{\mathrm{m}}\sim 10^{7}$ torsion micropendula based on mass-loaded Si3N4 nanoribbons, which achieve $S_{a}^{\mathrm{th}}\sim 10^{-9}\;g_{0}/\sqrt{\mathrm{Hz}}$ at room temperature [47]. Cryogenically cooled, these devices could reach the desired sensitivity for novel UDM searches in the 10 Hz to 1 kHz band, in conjunction with quantum-limited optical lever measurements [48, 49]; they also provide a route to vibration isolation through heterogeneous mass-loading [50], the latter of which has been demonstrated using a flip chip approach [46].
Finally, we highlight a recent proposal to search for UDM with an array of $N=10$ cryogenic membrane accelerometers probed by a distributed-squeezed light source [16], building on an $N=2$ demonstration with two $0.1$-$m$ membranes [51]. This approach enables enhanced sensitivity (as much as $N$-fold [16]) if the power per sensor is fixed and, therefore, a route towards improved UDM searches with a cryogenic detector limited by photothermal heating or other power constraints. Loss and scalability are key challenges. Towards this end, an emerging class of photonic-integrated accelerometers [52] and squeezed light sources [53] offer a promising route.
[…]
I Acknowledgments
The authors thank Swati Singh and Daniel Grin for laying the theoretical foundations of this work, Atkin Hyatt for designing and fabricating next generation test masses presented in the Supporting Materials, Christian Pluchar for guidance with laser intensity noise reduction, Andrew Land for help constructing the vibration isolation system, and Utkal Pandurangi for helpful discussions. This work was supported by NSF Grants 2209473 and 1945832 and from the Northwestern University Center for Fundamental Physics and the John Templeton Foundation through a Fundamental Physics grant.